The slope of the line whose equation is 3y + 2x = 1 is?


Answer by nerdybill(7384)

Put:
3y + 2x = 1
into "slope-intercept" form of y=mx+b
.
3y + 2x = 1
2x = -3x+1
x = (-3/2)x + 1/2
.
From inspection, we see that the slope is -3/2